OpenStudy (xapproachesinfinity):

I'm trying to prove this : \[\int \sin^n x\cos^m x dx=\frac{-\sin^{n-1}x\cos^{m+1}x}{n+m}+\frac{n-1}{n+m}\int \sin^{n-2}x \cos^mx dx\]

OpenStudy (xapproachesinfinity):

this is avery interesting identities for this kind of integrals

OpenStudy (adamaero):

mathematica.stackexchange.com Better for proofs

OpenStudy (xapproachesinfinity):

i was trying with \[u=\sin^{n-1}x\cos^mx\] \[dv=\sin x\]

OpenStudy (xapproachesinfinity):

but it was very messy calculations lol

OpenStudy (xapproachesinfinity):

i have also thought of doing something of this sort \[\sin^nx=\sin^2x\sin^{n-2}x \]

OpenStudy (xapproachesinfinity):

i thought of making it somewhat easy to integrate problem but i hit the wall lol

OpenStudy (xapproachesinfinity):

@jim_thompson5910

OpenStudy (xapproachesinfinity):

the last cos should be power m not n

OpenStudy (freckles):

so you have tried to do integration by parts?

OpenStudy (xapproachesinfinity):

yes that is my attempt

OpenStudy (xapproachesinfinity):

was*

OpenStudy (xapproachesinfinity):

'm lagging did you write something?

OpenStudy (freckles):

\[\sin^{n-1}(x) \sin(x) \cos^m(x)=\sin^n(x)\cos^m(x) \\ \int\limits_{}^{} \sin^{n-1}(x) \cdot \cos^m(x) \sin(x) dx \\ \] integrate the cos^m(x)*sin(x) and derivative the sin^(n-1)(x) let's see what happens ...

OpenStudy (xapproachesinfinity):

hmm let me try that i did the opposite i chose dv=sinx instead of sin^(n-1)x

OpenStudy (freckles):

\[I=\int\limits \sin^{n-1}(x) \cdot \cos^m(x) \sin(x) dx \\ I=\sin^{n-1}(x) \cdot \frac{-\cos^{m+1}(x)}{m+1} - \int\limits (n-1)\sin^{n-2}(x) \cos(x) \cdot \frac{-\cos^{m+1}(x)}{m+1} dx\] \[I=\frac{-1}{m+1} \sin^{n-1}(x) \cos^{m+1}(x)+\frac{n-1}{m+1}\int \sin^{n-2}(x) \cos^{m+2} dx\] is this what you got?

OpenStudy (freckles):

for the first round

OpenStudy (xapproachesinfinity):

I'm still in the process i should have something let me see

OpenStudy (freckles):

I missed the (x) thingy at the end of my line

OpenStudy (xapproachesinfinity):

we are in the same page

OpenStudy (xapproachesinfinity):

don't worry about it i got you:)

OpenStudy (freckles):

I guess we can see what happens if we apply integration by parts again :p

OpenStudy (xapproachesinfinity):

oh i got it

OpenStudy (freckles):

you got it?

OpenStudy (xapproachesinfinity):

i mean i can see it from here :) i was stuck on using the right u and dv

OpenStudy (freckles):

coolness :)

OpenStudy (xapproachesinfinity):

sometimes i touch around the problem i get close and get shy haha

OpenStudy (xapproachesinfinity):

like here i was on the right road but didn't manage it

OpenStudy (freckles):

because it could mean a lot of work and no happiness :(

OpenStudy (xapproachesinfinity):

yes indeed!

OpenStudy (xapproachesinfinity):

you hit the answer right way that's quite a skill you have:)

OpenStudy (xapproachesinfinity):

away*

OpenStudy (xapproachesinfinity):

i have another intersting integral i want to challenge it and see what can i come up with \[\int_{0}^{\pi/2}\sin^nxdx= \] don't even remember the rest lol some polynomial

OpenStudy (xapproachesinfinity):

i manged to see that part actually that is why i made sinx sin^(n-1)x cos^mx too

OpenStudy (freckles):

I keep mistyping what I mean so I gave up

OpenStudy (freckles):

polynnomial?

OpenStudy (xapproachesinfinity):

hahaha no worries :) thanks a lot, will try to tackle that problem later i have 3 more to do all of them have to do with by parts

OpenStudy (xapproachesinfinity):

yeah i have to prove it can be written as a polynomial or something i don't remember the rest. i need to look at the notes i think the problem are taken from calculus sullivan

OpenStudy (freckles):

n is natural right

OpenStudy (xapproachesinfinity):

yes sure it is about cases n odd n even they give different results

OpenStudy (xapproachesinfinity):

then part b asks about polynomial something

OpenStudy (freckles):

right even n gives us irrational number (the irrational part being the pi thing) and the odd n gives us rational number I was trying to find a pattern between the answers (well looking at the evens and odds separate)

OpenStudy (xapproachesinfinity):

yeah it is about that

OpenStudy (freckles):

I want to go ahead and see if we can look at the odd n in general let n=2k+1 where k>0 is an integer \[\int\limits_{0}^{\frac{\pi}{2}}\sin^{2k+1}(x) dx \\ \int\limits_{0}^{\frac{\pi}{2}} \sin(x) \sin^{2k}(x) dx \\ \int\limits_0^{\frac{\pi}{2}} \sin(x)(1-\cos^2(x))^k dx\]

OpenStudy (freckles):

here we can do a sub

OpenStudy (xapproachesinfinity):

yes seems u=cosx

OpenStudy (freckles):

\[u=\cos(x) \\ du=-\sin(x) dx \\ -\int\limits_{1}^{0}(1-u^2)^k du \\ \int\limits_0^1(1-u^2)^k du\]

OpenStudy (freckles):

thinking...

OpenStudy (xapproachesinfinity):

hmm good

OpenStudy (freckles):

http://openstudy.com/users/freckles#/updates/54de8a16e4b0b0ad8854cba3 we did something really cool here that I'm thinking maybe we can do something similar here maybe

OpenStudy (freckles):

But.. I have bad news

OpenStudy (freckles):

I have to retire tonight I shall be back tomorrow hopefully

OpenStudy (xapproachesinfinity):

eh i'm leaving too i will try myself to see if there is anything i can come up with those triple integral make me dizzy hahaha

OpenStudy (freckles):

I was talking about the result we derived there \[I(m,n) =\int\limits_0^a x^m(a-x)^n = \dfrac{m!n!}{(m+n+1)!}a^{m+n+1}\] This result was obtained by integration by parts

OpenStudy (xapproachesinfinity):

oh i see

OpenStudy (freckles):

We will of course not be able to use that result because our integrand is a bit different

OpenStudy (xapproachesinfinity):

but the idea is related to that correct

OpenStudy (freckles):

but we might be able to do integration by parts to obtain some kinda output like that above

OpenStudy (xapproachesinfinity):

oh i see

OpenStudy (freckles):

I could be wrong I haven't tried it

OpenStudy (freckles):

but i will tomorrow

OpenStudy (xapproachesinfinity):

well it is worth trying anyway :)

OpenStudy (freckles):

\[I=\int\limits_{0}^{1}(1-u^2)^k du \\ I= \int\limits_0^11 \cdot (1-u^2)^k du \\ I_1=u(1-u^2)^k|_0^1-\int\limits_0^1 u k (1-u^2)^{k-1} (-2u) du \\ I_1=\int\limits_0^1 k(1-u^2)^{k-1} 2u^2 du \\ I_2=\frac{2 u^3}{3} k (1-u^2)^{k-1}|_0^1-\int\limits_0^1 \frac{2u^3}{3}k(k-1)(1-u^2)^{k-2}(-2u) du \\ I_2=\int\limits_0^1 \frac{4 u^4}{3}k(k-1)(1-u^2)^{k-2} du \] We need to do this k times for I_k we should have (1-u^2)^(k-k) which means we will able to integrate other little factor u^some power . We can stop when we see the pattern: \[I_3=\frac{4 u^5}{ 3 \cdot 5 }k(k-1)(1-u^2)^{k-2}|_0^1-\int\limits_0^1 \frac{4 u^5}{3 \cdot 5 } k(k-1)(k-2)(1-u^2)^{k-3}(-2u) du \\ I_3=\int\limits_0^1 \frac{8 u^{6}}{3 \cdot 5} k(k-1)(k-2)(1-u^2)^{k-3} du\] .... hmm.. so the pattern looks sorta like this: \[I_n=\int\limits_0^1 \frac{2^n u^{2n}}{1 \cdot 3 \cdot 5 \cdot 7\cdots (2n-1)}[k(k-1)(k-2) \cdots (k-n)](1-u^2)^{k-n} du \] replace n with k since we wanted k times \[I_k=\int\limits_0^1 \frac{2^k u^{2k}}{\Pi_{i=1}^{k}(2i-1)} k! (1-u^2)^{k-k} du \\ I_k=\int\limits_0^1 \frac{2 ^k u^{2k}}{\Pi_{i=1}^{k} (2i-1)} k! du =\frac{2^k k!}{\Pi_{i=1}^{k}(2i-1)} \cdot \frac{u^{2k+1}}{2k+1}|_0^1 \\ =\frac{2^k k!}{(2k+1) \cdot \Pi_{i=1}^{k}(2i-1)}\] We might want to check that and see if I made any error .

OpenStudy (freckles):

so that means we have \[I_n=\int\limits_{0}^{\pi/2}\sin^nxdx \\ I_{2n+1}=\frac{2^k k!}{(2k+1) \cdot \Pi_{i=1}^{k}(2i-1)}\]

OpenStudy (freckles):

Now we need to think of the even case

OpenStudy (freckles):

and honestly I think about this case a little more

OpenStudy (freckles):

type-o in my one line: \[I_n=\int\limits_0^1 \frac{2^n u^{2n}}{1 \cdot 3 \cdot 5 \cdot 7\cdots (2n-1)}[k(k-1)(k-2) \cdots (k-(n-1))](1-u^2)^{k-n} du\]

OpenStudy (xapproachesinfinity):

thanks @freckles i was pretty close to that pattern you found too

OpenStudy (freckles):

I can't find a way to get something like that for the even n route

OpenStudy (freckles):

@ganeshie8 Is it possible to get a solution similar to that the even route way ?

ganeshie8 (ganeshie8):

@freckles all the definite integrals in this thread are just a special case of \[I(m,n) =\int\limits_0^a x^m(a-x)^n = \dfrac{m!n!}{(m+n+1)!}a^{m+n+1}\]

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{1}(1-u^2)^k du \] let \(u^2 = t \implies du = \dfrac{t^{-1/2}dt}{2}\) the integral becomes \[\frac{1}{2}\int\limits_0^1 t^{-1/2} (1-t)^k dt = \dfrac{(-1/2)! k!}{2(-1/2+k+1)!}= \dfrac{k!}{2(k+1/2)!}\sqrt{\pi}\] this is probably not what the assignment is about.. this form certainly looks a lot compact however xD

ganeshie8 (ganeshie8):

also we dont need to consider cases (n odd, even) if we allow rational and negative factorials in answer..

ganeshie8 (ganeshie8):

Let \(n\) be any integer : \[\int\limits_{0}^{\pi/2}\sin^nxdx= \int\limits_{0}^{\pi/2}(\sin^2 x)^{(n-1)/2} \sin xdx = \frac{1}{2}\int\limits_0^1 u^{-1/2}(1-u)^{(n-1)/2} du \\~\\~\\= \dfrac{(\frac{-1}{2})! (\frac{n-1}{2})!}{2(\frac{-1}{2}+\frac{n-1}{2}+1)!} = \dfrac{ (\frac{n-1}{2})!}{2(\frac{n}{2})!} \sqrt{\pi}\]

OpenStudy (xapproachesinfinity):

what about deriving \[\int_{0}^{\pi/2}\sin^nxdx=\frac{(n-1)(n-3)......(4)(2)}{n(n-2)...........(5)(3)(1)}\] for n odd

ganeshie8 (ganeshie8):

we use the main formula : \[\int \sin^n x\cos^m x dx=\frac{-\sin^{n-1}x\cos^{m+1}x}{n+m}+\frac{n-1}{n+m}\int \sin^{n-2}x \cos^mx dx\] set \(m=0\) and plugin the bounds

OpenStudy (xapproachesinfinity):

@ganeshie8 you mean for my last comment i should use that to derive that fraction

ganeshie8 (ganeshie8):

that will do right ? why derive everything again..

OpenStudy (xapproachesinfinity):

because the first question is to derive that massive fraction